[Ncep.nhc.nco_contacts] TSB Morning Rounds - Sun Mar 13, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Mar 14 13:20:46 UTC 2016


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
TAFB and APO are still working through resolving the GFE missing grid  
banners
HSU:
None
Other items:
1. AWIPS2 Collaboration test will take place today from 3-4pm EDT
2. NCO is still investigating the GOES-W dropouts that were reported last  
week. Satellite processing will continue to run from the NESDIS feed until  
that is resolved.
3. NCO is still working on restoring ssh from the NHCN LDAD to all the  
webservers

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160314/8aa847a3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list